Mrs.S invests a part of Rs.144,000 in 12% stock having face value of Rs.100 at a price of Rs.120 and the remainder in 15% stock having face value of Rs.100 at a price of Rs.125. If his total dividend per annum is Rs.16,320 how much does she invest in 12% stock at Rs.120??
1
Expert's answer
2012-05-29T08:26:55-0400
Total dividend per annum: 16320 = (100 / 120) * x * 0.12 + (100 / 125)
*
(144000 - x) * 0.15. So,
16320 = x * 0.1 + 17280 - 0.12 * x
0.02 * x =
17280 - 16320
x = 960/ 0.02 = 48000 invested in 12% stock at Rs.120